### v2

@@ 1 Background
In order to address accuracy loss in the presence of sequencing errors, strobemers [19] were introduced to concatenate selected minimizers to improve indexing and error tolerance. Alternatively, syncmers [20] select k-mers based on specific subsequence properties, offering better conservation, lower sampling density, and higher error tolerance compared to classical minimizers.
As an alternative to the aforementioned sketching techniques, Locally Consistent Parsing (i.e., LCP) [21, 22] proposes a string partitioning method with consistent data reduction for building dictionaries like indexes. Sahinalp and Vishkin introduced LCP based on the Deterministic Coin Tossing (DCT) [23] technique, to construct suffix trees in parallel efficiently. LCP was also applied to pattern matching [21, 24] and edit distance computation [25]. A practical but non-iterative variant of LCP was implemented to boost the Lempel-Ziv algorithm for compressing biological sequences [26]. Later, locally consistent grammar-based text compression was proposed in  [27]. Other applications of LCP, such as approximate string comparison and text indexing, can be found in  [28, 29, 30].
In this article, we introduce a generalized, iterative, and practical implementation of LCP, called Lcptools, providing an experimental analysis of the properties of the cores compared to minimizers, syncmers, MinHash sketches and UHS. The analysis highlights that LCP cores are fewer in number and exhibit longer but more consistent distribution throughout the processed string. We also demonstrate the utility of LCP for variation graph construction using our novel algorithm LCPan that outperforms vg, the state-of-the-art variation graph constructor, in both running time ($10\times$) and memory usage ($13\times$), while demonstrating slightly better alignment accuracy using GraphAligner [31] on HiFi data. We note that both LCPan and vg generate variation graphs using a linear reference genome and a VCF file that contains genomic variation, where alternatives like pggb [32] and Minigraph-Cactus [33] build variation graphs using fully assembled contigs and scaffolds. We therefore limit our comparisons to only using vg for data consistency.
@@ 2 Results
Locally Consistent Parsing (LCP) is a string processing technique that partitions and labels strings into nearly equal-length substrings, known as cores [21, 22]. Unlike sketching techniques, LCP ensures that (1) the cores have a uniform positional distribution over the input string, (2) identical cores share the same labels, and (3) each character in the input string is included in at least one core.
@@ 2.1 Properties of LCP Cores
We then compared the LCP method against commonly used sketching methods, namely minimizers, syncmers, UHS, and MinHash, using the same input genome to assess the performance of each method regarding the number of selected k-mers and the distances between each selected k-mer. The LCP levels were selected as 2 and 3 because the average length of the cores is similar to the commonly used k-mer sizes in practice. We selected $k=15$ and $w=10$ for the canonical minimizers as these are the default values in the popular minimizer-based alignment tool, Minimap2 [16]. For the syncmer analysis, we selected $k=15$, $s=4$, and $t=0$, allowing s-mers to be located only at the beginning of the k-mer, based on the recommended settings in the original publication [20]. For the UHS analysis, we downloaded the pre-computed k-mer sets released by the authors [17] [^1] https://acgt.cs.tau.ac.il/docks/ for $k=13$ and $L=20$. We then mapped these k-mers to the CHM13v2.0 assembly and retained only those that matched exactly. Finally, for MinHash, we analyzed the default Mash sketch size of 1,000 k-mers, and additionally 5 million k-mers ($k=31$), which roughly corresponds to the number of cores in LCP level 8.
@@
As shown in Table 2, for $k=15$, we found over 665 million minimizers in CHM13v2.0, among which 92.6 million were unique. Syncmer count was smaller, at 371 million with 58 million unique. Since UHS are computed independently from a reference genome, the unique count was smaller, at 20.6 million. Still, when mapped to the genome, it returned 983.5 million locations, the highest value among all techniques tested. This is likely because the UHS set contained many k-mers with long homopolymers. MinHash is the coarsest-grained sketching method in our tests. As expected, it showed the smallest number of k-mers, which is a user-defined parameter. Similar to our LCP analysis, we calculated the average distances between each pair of minimizers, syncmers, UHS, and MinHash. On average, the distance between two consecutive minimizers was 4.69 nucleotides for $k=15$, whereas the distance between two consecutive syncmers was 8.40 nucleotides. UHS k-mers were more densely organized in the genome, where the average distance was 3.16 nucleotides. Once again, MinHash demonstrated the widest range of distances between selected k-mers, with 454.42 bp average distance for the large sketch size (5 million) and 2.78 Mb distance for the default sketch size of 1,000 in Mash. The cores in LCP level-2 were, on average, 10.47 bp, and in level-3, were, on average, 26.58 bp; the former is slightly shorter, and the latter is slightly longer than the minimizer and syncmer lengths we considered. Although shorter, the number of level-2 cores was less than that of minimizers for $k=15$. Importantly, the number of unique level-2 cores is several orders of magnitude smaller than the minimizers or syncmers of any size, thanks to consistent identification of cores using LCP. The number of level-3 cores is less than all the alternatives we considered, except MinHash. Perhaps most importantly, among all the other sketches, LCP deviates less proportionally from the average distances. Higher LCP levels show even lower numbers of cores, therefore reducing the computational cost for storage, indexing, and query (Table 1).
@@
We note that finding minimizers and syncmers requires only a linear scan of the input string and simple lexicographic comparison, therefore faster than identifying cores in LCP in practice. Computing UHS, on the other hand, is NP-hard [17]; thus, the UHS k-mer sets can only be approximated. However, the granularity of sketches generated by LCP is adjustable using different levels. Therefore, depending on the specific use case, longer but fewer or shorter but more frequent cores can be used. On the other hand, higher LCP levels are constructed by parsing lower levels; therefore, cores at several levels can be computed in one pass of the input string. Similar flexibility by minimizers and syncmers can be achieved only by changing the parameters $k$, $w$, and $s$, which will require multiple passes on the input string and incur significant bottlenecks on both the sketching method’s run time and memory footprint. However, we leave a rigorous comparison of LCP with sketching methods for read mapping as future work.
@@ 2.2 Properties of LCPan Graphs and Construction Performance
  • We compared graph construction metrics between LCPan and VG using the HPRC data set [35] on GRCh38. The table shows the total number of segments, total number of links, execution time, average segment length, standard deviation of segment length, memory usage (RSS), and output size for different LCP levels of LCPan and VG.
@@ 2.3 Scaling to multiple threads
Figure 1: Multi-thread scaling analysis of LCPan using HPRC data on GRCh38. Memory and run time of graph construction using different numbers of threads.
LCPan first processes the reference genome to build the backbone graph sequentially and then assigns the set of variations to multiple threads to amend the graph. The partition consistency provided by LCP enables LCPan to efficiently scale computation into multiple threads. To evaluate the parallelization performance, we constructed the same variation graph using different numbers of threads (Figure 1). We observed that the memory usage was not significantly affected across different thread settings. Meanwhile, LCPan speed was improved up to 16 threads, and more threads did not improve the run time. This plateau in thread scaling is caused by two limiting factors: sequential processing of the reference genome and the task allocation speed of the main thread that distributes the processing to the worker threads.
@@ 2.4 Alignment accuracy
We constructed both LCPan and vg variation graphs using only chromosomes 1, 10, and 22 (GRCh38) and the variants released by the Human Pangenome Reference Consortium (HPRC) [35]. vg constructed the graph in 6m 18s using 12.91 GB of memory, while LCPan constructed the graph in only 2m 16s using 2.17 GB of memory. We set the k-mer size in the vg graph to 64 and the LCP level to 5 for LCPan to make the average segment sizes (i.e., lengths of the sequences represented in nodes) comparable. The resulting graphs included segments of average lengths 56.25 bp and 55.17 bp for LCPan and vg, respectively.
@@
Table 4: GraphAligner alignment results for PacBio HiFi and ONT reads aligned on LCPan and VG graphs from chromosomes 1, 10, and 22.
@@
Dataset PacBio HiFi ONT PacBio HiFi ONT
Input Reads 820,665 281,312 820,665 281,312
Reads w/ align. 820,665 281,306 820,665 281,308
Align. Time ($h$:$m$:$s$) 06:49:06 06:19:06 08:12:07 07:42:47
Memory (GB) 82.01 114.57 83.08 103.29
Precision 0.40 0.41 0.38 0.41
Recall 0.70 0.70 0.72 0.71
F1 0.51 0.51 0.50 0.52
@@ 4 Methods
Below, we describe a simple variant of LCP, expressed in terms of four rules, to identify cores in an input string from a given alphabet (e.g., the 44-letter DNA alphabet). We will later show how to apply LCP iteratively using DCT in Section 4.1.
@@
    Stranded Sequence Rule (SSEQ): A substring $w,|w|>2$ is a core if its characters are either strictly increasing or decreasing w.r.t. the lexicographic order, and only the first and last characters are part of either a LMIN, LMAX, or a RINT cores. Specifically, if $w=xyz{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{0}\dots a_{n}}klm$, where $n\geq 1$ and xyzxyz and klmklm are identified as cores and $z<a_{0}<\dots<a_{n}<k$ or $z>a_{0}>\dots>a_{n}>k$, then $z{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{0}\dots a_{n}}k$ is a core (SSEQ type).
Figure 2 depicts an example of the cores identified in a short DNA sequence. As the cores are identified, the two-bit encoding of the underlying characters (i.e., A = $0$, C = 01, G = 10, T = 11) is used to form the core’s bitstream. Note that we use the term core alphabet to refer to the bitstreams that form the cores, specifically the DNA encoding used in the first LCP iteration.
@@
Adjacency Property: within a substring of length 33, at most 22 characters can be the starting positions of the cores.
@@ 4.1 Iterative Application of Locally Consistent Parsing
Our goal is to apply LCP iteratively to increase the distances between consecutive cores while reducing their number and increasing their length by a constant factor. However, since the core bitstreams are larger than the string’s alphabet (in our context, DNA), directly applying the LCP rules to these bitstreams would increase the distances without constraints. To address this, starting from the second LCP iteration ($i>1$), the bitstreams and LCP operations depend on the reduction in alphabet size via DCT. [^6] When fully applied, the DCT reduces the alphabet size to 22, enabling a uniform bitstream length of 66, which is crucial for maintaining consistency in later LCP iterations. After this reduction in alphabet size, the string of core bitstreams will satisfy both Lemmas 1 and 2, maintaining an even distance distribution and compact length; thus, LCP can be applied iteratively.
@@
As described in Figure 3, DCT takes the bitstreams of cores at level $i$ as input, compares each core’s bitstream with its left neighbor, and finds the least significant bit where it differs from the right. Then, it replaces the original bitstream with a shorter bit string formed by concatenating the index of the differing bit with the value of the corresponding bit. If the number of bits representing a core bitstream is $k$, DCT will produce a reduced bitstream of length $\log(k)+1$. Here, $k$ represents the number of bits in the original core bitstream, and $\log(k)+1$ corresponds to the binary representation of the differing bit’s index and its value [^7] It is possible to apply DCT again to the DCT-reduced core bitstreams to further decrease the alphabet size. However, in this implementation, we apply DCT only once..
@@
Thus, in iteration $i>1$, after applying DCT on bitstreams and then LCP on newly computed alphabets, we can identify level-$(i+1)$ cores. Let $x$ be the core at level-$(i+1)$, which is composed of level-$i$ cores $y$, $z$, $w,\ldots$, i.e. $x=yzw\ldots$ Suppose that the DCT reduced alphabet size for these cores is $y^{\prime},z^{\prime},w^{\prime},\ldots$; then the level-$(i+1)$ core bitstream for $x$ will be the concatenation of these bit strings, namely, $y^{\prime}z^{\prime}w^{\prime}\ldots$. Since DCT consistently reduces each occurrence of the combination $yzw\ldots$ to the same bitstream representation, the cores are guaranteed to align consistently in the resulting bitstream.
Equipped with DCT, each iteration of LCP will reduce the number of cores while increasing the average length of a core and the average distance between two consecutive cores by a constant factor $c$. As shown earlier, while this constant $c$ is at least $3/2$, in practice $c\sim 2.34$ as demonstrated in Table 1. The cores identified in iteration $i$ are said to be of level-$i$ cores. There will be $\sim n/c^{i}$ cores of level-$i$, representing substrings of length $O(c^{i})$; the average distance between consecutive cores will be $c^{i}$.
LCP runs in time linear with the length of the input string at each level. Since the number of cores decreases geometrically by a factor of $c$ at each level, the running time of LCP for all levels is linear in the length of the original input string.

### v3

@@ 1 Background
In order to address accuracy loss in the presence of sequencing errors, strobemers [19] were introduced to concatenate selected minimizers, thereby improving indexing and error tolerance. Alternatively, syncmers [20] select k-mers based on specific subsequence properties, offering better conservation, lower sampling density, and higher error tolerance compared to classical minimizers.
As an alternative to the aforementioned sketching techniques, Locally Consistent Parsing (i.e., LCP) [21, 22] proposes a string partitioning method with consistent data reduction for building dictionaries like indexes. Sahinalp and Vishkin introduced LCP based on the Deterministic Coin Tossing (DCT) [23] technique, to construct suffix trees in parallel efficiently. LCP was also applied to pattern matching [21, 24] and edit distance computation [25]. A practical but non-iterative variant of LCP was implemented to boost the Lempel-Ziv algorithm for compressing biological sequences [26]. Later, locally consistent grammar-based text compression was proposed in  [27]. Other applications of LCP, such as approximate string comparison and text indexing, can be found in  [28, 29, 30]. However, except for the non-iterative variant for sequence compression [26], none of the previously proposed algorithms aim to process biological sequences, and only two open-source implementations are currently available [27, 30]. We note that these implementations are specifically for text indexing and compression, and they do not provide support for utilizing LCP for various use cases.
Locally Consistent Parsing (LCP) is a string processing technique that partitions and labels strings into nearly equal-length substrings, known as cores [21, 22]. We note that LCP is not a sketching method; however, the cores can potentially serve as an alternative to sketching. Unlike sketching techniques, LCP ensures that (1) the cores have a uniform positional distribution over the input string, (2) identical cores share the same labels, and (3) each character in the input string is included in at least one core. In this article, we introduce a generalized, iterative, and practical implementation of LCP, called Lcptools, providing an experimental analysis of the properties of the cores compared to minimizers, syncmers, MinHash sketches and UHS. Our analysis highlights that LCP cores are fewer in number and exhibit longer but more consistent distribution throughout the processed string. We also demonstrate the utility of LCP for variation graph construction using our novel algorithm LCPan that outperforms vg, the state-of-the-art variation graph constructor, in both running time ($10\times$) and memory usage ($13\times$), while demonstrating slightly better alignment accuracy using GraphAligner [31] on HiFi data. We note that both LCPan and vg generate variation graphs using a linear reference genome and a VCF file that contains genomic variation, where alternatives like pggb [32] and Minigraph-Cactus [33] build variation graphs using fully assembled contigs and scaffolds. We therefore limit our comparisons to only using vg for data consistency.
@@ 2.1 Properties of LCP Cores
This observation empirically supports the theoretical analysis that $2\leq c\leq 3$, as originally developed by Sahinalp and Vishkin [22]. We note that no k-mer-based sketching scheme (e.g., minimizers and syncmers) can guarantee fixed distances between consecutive substrings. Minimizers are only guaranteed to have a maximal gap of size $w$ between them. We refer the reader to other work in the literature for an analysis of minimizer densities [35].
@@
We then compared the LCP method against commonly used sketching methods, namely minimizers, syncmers, UHS, and MinHash, using the same input genome to assess the performance of each method regarding the number of selected k-mers and the distances between each selected k-mer. The LCP levels were selected as 2 and 3 because the average length of the cores is similar to the commonly used k-mer sizes in practice. We selected $k=15$ and $w=10$ for the canonical minimizers as these are the default values in the popular minimizer-based alignment tool, Minimap2 [16]. For the syncmer analysis, we selected $k=15$, $s=4$, and $t=0$, allowing s-mers to be located only at the beginning of the k-mer, based on the recommended settings in the original publication [20]. For the UHS analysis, we downloaded the pre-computed k-mer sets released by the authors for k=13 and L=20. Similarly, for MinHash, we analyzed the default Mash sketch size of 1,000 k-mers, as well as 5 million k-mers (k=31), which roughly corresponds to the number of cores in LCP level 8. We then mapped these k-mers to the CHM13v2.0 assembly and retained only those that matched exactly. Note that because of the repetitions in the genome, the number of observed k-mers in the genome is slightly higher than the k-mer counts in UHS and MinHash sketches.
@@
As shown in Table 2, for $k=15$, we found over 665 million minimizers in CHM13v2.0, among which 92.6 million were unique. The syncmer count was smaller, at 371 million, with 58 million unique users. Since UHS are computed independently from a reference genome, the unique count was smaller, at 20.6 million. Still, when mapped to the genome, it returned 983.5 million locations, the highest value among all techniques tested. This is likely because the UHS set contained many k-mers with long homopolymers. MinHash is the coarsest-grained sketching method in our tests. As expected, it showed the smallest number of k-mers, which is a user-defined parameter. Similar to our LCP analysis, we calculated the average distances between each pair of minimizers, syncmers, UHS, and MinHash. On average, the distance between two consecutive minimizers was 4.69 nucleotides for $k=15$, whereas the distance between two consecutive syncmers was 8.40 nucleotides. UHS k-mers were more densely organized in the genome, where the average distance was 3.16 nucleotides. Once again, MinHash demonstrated the widest range of distances between selected k-mers, with 454.42 bp average distance for the large sketch size (5 million) and 2.78 Mb distance for the default sketch size of 1,000 in Mash. The cores in LCP level-2 were, on average, 10.47 bp, and in level-3, were, on average, 26.58 bp; the former is slightly shorter, and the latter is slightly longer than the minimizer and syncmer lengths we considered. Although shorter, the number of level-2 cores was less than that of minimizers for $k=15$. Importantly, the number of unique level-2 cores is several orders of magnitude smaller than the minimizers or syncmers of any size, thanks to consistent identification of cores using LCP. The number of level-3 cores is less than all the alternatives we considered, except MinHash. Perhaps most importantly, among all the other sketches, LCP deviates less proportionally from the average distances. Higher LCP levels show even lower numbers of cores, therefore reducing the computational cost for storage, indexing, and query (Table 1).
@@
We note that finding minimizers and syncmers requires only a linear scan of the input string and simple lexicographic comparison, therefore faster than identifying cores in LCP in practice. Computing a minimum-size UHS, on the other hand, is NP-hard [17]; thus, the UHS k-mer sets can only be approximated. However, the granularity of sketches generated by LCP is adjustable using different levels. Therefore, depending on the specific use case, longer but fewer or shorter but more frequent cores can be used. On the other hand, higher LCP levels are constructed by parsing lower levels; therefore, cores at several levels can be computed in one pass of the input string. Similar flexibility by minimizers and syncmers can be achieved only by changing the parameters $k$, $w$, and $s$, which will require multiple passes on the input string and incur significant bottlenecks on both the sketching method’s run time and memory footprint. However, we leave a rigorous comparison of LCP with sketching methods for read mapping as future work.
@@ 2.2 Properties of LCPan Graphs and Construction Performance
  • We compared graph construction metrics between LCPan and VG using the HPRC data set [36] on GRCh38. The table shows the total number of segments, total number of links, execution time, average segment length, standard deviation of segment length, memory usage (RSS), and output size for different LCP levels of LCPan and VG. Graphs are constructed using all chromosomes.
Table 4: Yeast pangenome construction using 100 strains with LCPan, VG, and VariantStore
Metric LCPan (5) VG VariantStore
Total # Segments 252,142 437,196 73,080
Total # Links 276,486 461,631 97,440
Avg. Seg. Len. (bp) 47.97 27.67 165.51
StdDev Seg. Len. (bp) 31.86 10.34 10.87
Exec. Time (sec) 0.52 2.25 9.61
Peak memory (MB) 34.63 84.17 100.80
Output r/GFA (MB) 14 20 N/A
  • Comparison of graph construction metrics between LCPan (level 5), VG, and VariantStore using 100 yeast strains [37] on S. cerevisiae assembly R64 (GCA_000146045). VariantStore values represent the overall values for all chromosomes.
[…]
@@ 2.3 Scaling to multiple threads
Figure 1: Multi-thread scaling analysis of LCPan and vg using HPRC data on GRCh38. Memory and run time of graph construction using different numbers of threads. LCPan and vg were run using their native multithreading implementations. vg was additionally parallelized using GNU Parallel by distributing single-threaded processes on chunks across concurrent jobs (VG-GNU).
LCPan first processes the reference genome to build the backbone graph sequentially and then assigns the set of variations to multiple threads to amend the graph. The partition consistency provided by LCP enables LCPan to efficiently scale computation into multiple threads ( see Section 4.4 for details). To evaluate the parallelization performance, we constructed the same variation graph using different numbers of threads for both LCPan and vg (Figure 1). We note that vg documentation recommends parallel construction in two ways: 1) a command-line parameter (vg construct -t), and 2) by partitioning the genome and VCF file into chunks and distributing the task across multiple threads using the GNU Parallel tool [39]. We have used both methods for vg in our evaluation.
We observed that the memory usage was not significantly affected across different thread settings for both LCPan and vg. Meanwhile, LCPan speed was improved up to 8 threads, and more threads did not improve the run time. This plateau in thread scaling is caused by two limiting factors: the sequential processing of the reference genome and the task allocation speed of the main thread, which distributes the processing to the worker threads.
For vg, we observed that the native command-line parameter had no effect on improving efficiency, whereas the GNU Parallel-based technique provided speedups of up to 16 threads. In comparison, LCPan exhibits better scalability in both cases.
@@ 2.4 Alignment accuracy
We first constructed both LCPan and vg variation graphs using only chromosomes 1, 10, and 22 (GRCh38) and the variants released by the Human Pangenome Reference Consortium (HPRC) [36]. vg constructed the graph in 6m 18s using 12.91 GB of memory, while LCPan constructed the graph in only 2m 16s using 2.17 GB of memory. We set the k-mer size in the vg graph to 64 and the LCP level to 5 for LCPan to make the average segment sizes (i.e., lengths of the sequences represented in nodes) comparable. The resulting graphs included segments of average lengths 56.25 bp and 55.17 bp for LCPan and vg, respectively.
@@
Table 5: GraphAligner alignment results for PacBio HiFi and ONT reads aligned on LCPan and VG graphs.
@@
Dataset PacBio HiFi PacBio HiFi ONT PacBio HiFi PacBio HiFi ONT
Input Reads 820,665 5,764,123 281,312 820,665 5,764,123 281,312
Reads w/ align. 820,665 5,764,122 281,306 820,665 5,764,123 281,308
Align. Time (hh:mm:ss) 06:49:06 63:46:26 06:19:06 08:12:07 91:34:26 07:42:47
Memory (GB) 82.01 220.08 114.57 83.08 411.91 103.29
Precision 0.40 0.49 0.41 0.38 0.38 0.41
Recall 0.70 0.53 0.70 0.72 0.72 0.71
F1 0.51 0.51 0.52 0.50 0.50 0.52
[…]
@@ 4 Methods
Below, we describe a simple variant of LCP, expressed in terms of four rules, to identify cores in an input string from a given alphabet (e.g., the $4$-letter DNA alphabet). We will later show how to apply LCP iteratively using DCT in Section 4.1.
@@
    Stranded Sequence Rule (SSEQ): A substring $w,|w|>2$ is a core if its characters are either strictly increasing or decreasing w.r.t. the lexicographic order, and only the first and last characters are part of either a LMIN, LMAX, or a RINT cores. Specifically, if $w=xyz{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{1}\dots a_{n}}klm$, where $n\geq 1$ and xyzxyz and klmklm are identified as cores and $z<a_{1}<\dots<a_{n}<k$ or $z>a_{1}>\dots>a_{n}>k$, then $z{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{1}\dots a_{n}}k$ is a core (SSEQ type).
Figure 2 depicts an example of the cores identified in a short DNA sequence. As the cores are identified, the two-bit encoding of the underlying characters (i.e., A = 00, C = 01, G = 10, T = 11) is used to form the core’s bitstream. Note that we use the term core alphabet to refer to the bitstreams that form the cores, specifically the DNA encoding used in the first LCP iteration.
@@
Adjacency Property: within a substring of length $3$, at most $2$ characters can be the starting positions of the cores.
@@ 4.1 Iterative Application of Locally Consistent Parsing
Our goal is to apply LCP iteratively, so that at higher levels, cores become longer, the distance between consecutive core start positions increases, and the total number of cores decreases, all by approximately a constant factor per level. A direct attempt to do this by reapplying the LCP rules to the raw bitstream representations of level‑$i$ cores is problematic: once we move beyond the DNA alphabet, the “symbols” (core bitstreams) effectively come from a much larger, variable-length alphabet, and applying the LCP core-identification rules on such an alphabet can lead to uncontrolled growth in spacing and loss of the even spacing properties ensured by the LCP lemmas at the base level.
@@
To keep iterative parsing well-behaved, we incorporate Deterministic Coin Tossing (DCT) when constructing levels $i>1$. DCT maps each level‑$i$ core bitstream to a shorter code over a reduced alphabet by comparing it to its immediate left neighbor (Figure 3). Concretely, let $b_{j-1}$ and $b_{j}$ be the (right-aligned) bitstreams of two consecutive level‑$i$ cores. DCT finds the least significant position $t$ (counting from the right, starting at 0) where $b_{j-1}$ and $b_{j}$ differ, and replaces $b_{j}$ by a new code that concatenates (i) the binary representation of $t$ and (ii) the value of $b_{j}$ at position $t$. If the original bitstreams have length $k$, then $t\in{0,\ldots,k-1}$ and the DCT code has length $\lceil\log k\rceil+1$ bits (the index plus one bit for the value). This reduction controls the effective alphabet size of the sequence on which LCP is applied, so that the resulting “string of reduced symbols” continues to satisfy the contiguity/adjacency-style constraints needed for iterative LCP.
[…]
@@
After DCT, we apply the standard LCP core rules to the resulting reduced-symbol sequence to obtain level‑($i+1$) cores. If a level‑($i+1$) core $x$ is the concatenation of consecutive level‑$i$ cores $y,z,w,\ldots$ (i.e., $x=yzw\ldots$), then the level‑($i+1$) bitstream for $x$ is defined as the concatenation of the corresponding DCT codes $y^{\prime},z^{\prime},w^{\prime},\ldots$. Because DCT is deterministic and depends only on adjacent core bitstreams, repeated occurrences of the same block of level‑$i$ cores yield the same reduced representation up to boundary effects, which is sufficient to support locally consistent parsing at higher levels.
Equipped with DCT, each LCP level reduces the number of cores while increasing both the average core length and the average distance between consecutive cores by a constant factor $c$. From the adjacency bound at each level, $c$ is at least 3/2 in the worst case; empirically, on the human genome, we observe $c\approx 2.34$ (Table 1). Thus, at level-$i$, we expect on the order of $n/c^{i}$ cores representing substrings of average length $\Theta(c^{i})$, and the average distance between consecutive core start positions is $\Theta(c^{i})$. Each level can be computed in time linear in the length of the level’s representation, and because the representation shrinks geometrically across levels, the total time to compute all levels is linear in the original input length.
Our implementation applies only one round of DCT as a pragmatic tradeoff between LCP’s guarantees and the efficiency of the “anchor”. Each additional DCT round requires increasing the core-substring length (and thus overlap between consecutive cores) to preserve the key property that core-ness of a substring is context independent (a core remains a core wherever it occurs). This added overlap between consecutive cores increases redundancy and reduces sparsity. As we have demonstrated, a single round of DCT already yields a stable, well-behaved empirical distribution of distances between consecutive cores on the human genome, so further rounds offer limited benefit relative to their cost. In rare cases, a single round can create a short stranded region not covered by any core substring; however, empirically, such regions are highly infrequent and do not measurably affect the overall distribution of distances between consecutive cores.